# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v1

@@ Abstract
In orthogonal frequency division multiplexing (OFDM), accurate channel estimation is crucial. Classical signal processing based approaches, such as minimum mean-squared error (MMSE) estimation, often require second-order statistics that are difficult to obtain in practice. Recent deep neural networks based methods have been introduced to address this; yet they often suffer from high complexity. This paper proposes an Attention-aided MMSE (A-MMSE), a novel model-based DNN framework that learns the optimal MMSE filter via the Attention Transformer. Once trained, the A-MMSE estimates the channel through a single linear operation for channel estimation, eliminating nonlinear activations during inference and thus reducing computational complexity. To enhance the learning efficiency of the A-MMSE, we develop a two-stage Attention encoder, designed to effectively capture the channel correlation structure. Additionally, a rank-adaptive extension of the proposed A-MMSE allows flexible trade-offs between complexity and channel estimation accuracy. Extensive simulations with 3GPP TDL channel models demonstrate that the proposed A-MMSE consistently outperforms other baseline methods in terms of normalized MSE across a wide range of SNR conditions. In particular, the A-MMSE and its rank-adaptive extension establish a new frontier in the performance complexity trade-off, redefining the standard for practical channel estimation methods.
Index Terms:
Orthogonal frequency division multiplexing (OFDM), Channel estimation, Minimum mean-squared error (MMSE), Attention Transformer, Deep learning
@@ I Introduction
Recently, inspired by the remarkable success of deep neural networks (DNNs) across various engineering fields, DNN-aided channel estimation for OFDM has gained considerable attention. DNNs, as pre-trained neural network models, can effectively capture complex data relationships without explicit prior statistical knowledge. Leveraging this data-driven approach, several prior works have explored DNN-aided OFDM channel estimation strategies. In [7], an OFDM channel slot was treated as image data. Based on this, a super-resolution network and a denoising neural network were simultaneously employed to estimate the OFDM channel (i.e., ChannelNet). In [8], using a similar modeling approach to [7], ReEsNet was proposed, where a residual learning-based DNN was used to reconstruct the OFDM channel. This method achieves lower computational cost and higher accuracy than [7]. In [9], a residual convolutional network combined with bilinear interpolation was proposed for OFDM channel estimation, offering enhanced flexibility to varying pilot patterns and reduced complexity. Beyond deterministic models, [10] optimized a deep generative model using compressive pilot measurements for high-dimensional channel estimation. This method leverages learned priors to reduce pilot overhead and surpass traditional sparse recovery techniques.
@@
Via numerical simulations across diverse 3GPP TDL scenarios, we show that the proposed A-MMSE and RA-A-MMSE significantly outperform traditional SP-based approaches (e.g., LS and MMSE) and recent DNN-based approaches (e.g., ChannelNet [7] and Channelformer [13]) in terms of channel estimation accuracy. In particular, the A-MMSE achieves approximately $43\%$ lower normalized MSE (NMSE) compared to conventional MMSE. We attribute these performance gains to the Attention Transformer’s ability to effectively capture nonstationary channel statistics, even under challenging high-mobility and frequency-selective fading conditions. Comparing to Channelformer, the A-MMSE achieves approximately $48\%$ lower NMSE. We note that these gain stem from our novel two-stage Attention encoder architecture, where frequency domain features are first extracted and subsequently refined through temporal correlation modeling. Moreover, the RA-A-MMSE further reduces computational complexity through adaptive rank reduction, maintaining over $99.6\%$ of the performance achieved by the full-rank A-MMSE. As a result, the RA-A-MMSE requires only $1.5\%$ of the computational complexity of Channelformer or MMSE (i.e., a $98.5\%$ reduction), while achieving approximately $36\%$ lower NMSE. The proposed methods also exhibit strong robustness to SNR mismatch, retaining nearly full performance ($\geq 98.1\%$) across a wide range of unseen SNR conditions. This underscores the practical efficiency and suitability of the proposed method for real-world 5G/6G implementations.
In summary, the proposed A-MMSE method offers several advantages over existing approaches. First, the proposed A-MMSE with novel two-stage Attention encoder effectively incorporates the OFDM channel correlation structure, thereby significantly improving channel estimation accuracy and robustness, achieving around $43\%$ and $48\%$ lower NMSE compared to MMSE and Channelformer, respectively. Second, once trained, the A-MMSE employs purely linear computations for OFDM channel estimation, significantly reducing computational complexity compared to conventional DNN-based methods that heavily rely on numerous nonlinear operations. Especially, the RA-A-MMSE dynamically adjusts its computational complexity according to available hardware resources, requiring $99.8\%$ fewer computations than the average of conventional DNN-based methods. This enables both efficient and flexible implementation in practical systems.
@@
Notations: We denote transpose by $(\cdot)^{\mathsf{T}}$, Hermitian transpose by $(\cdot)^{\mathsf{H}}$, complex conjugate by $(\cdot)^{*}$, column‑wise vectorization by $\operatorname{vec}(\cdot)$, real and imaginary parts of a complex quantity by $\Re(\cdot)$ and $\Im(\cdot)$, the magnitude by ||$\cdot$|| ⋅ |, and the Euclidean and Frobenius norms by $\lVert\cdot\rVert_{2}$ and $\lVert\cdot\rVert_{F}$, respectively. Element‑wise (Hadamard) multiplication is denoted by $\circ$, the Kronecker product by $\otimes$, and statistical expectation is written as $\mathbb{E}[\cdot]$. The set difference is denoted by $\setminus$, horizontal (row-wise) stacking by [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ], and vertical (column-wise) stacking by [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ].
@@ III-A SP-based Approach
The SP-based OFDM channel estimation strategies have been extensively studied in [6]. Among the various approaches including decision-directed channel estimation (DD-CE) and parametric model (PM)-based methods, this work intentionally focuses on pilot-assisted channel estimation, due to its robustness and practicality in wireless communication scenarios. In particular, the LS and MMSE estimators are commonly used as standard methods.
@@
The LS estimator requires no prior statistical knowledge of the channel or noise, making it simple and straightforward to implement. However, its performance may degrade significantly, especially at low signal-to-noise ratio (SNR) conditions or when the channel matrix ${\bf{H}}$ is ill-conditioned.
@@
Under the joint Gaussian assumption, the MMSE estimator is well-established to achieve optimal performance with respect to the MSE criterion. In particular, unlike the LS estimator, the MMSE estimator exploits knowledge of two covariance matrices: $\mathbf{R}_{\operatorname{vec}(\mathbf{H})\,\mathbf{H}_{p}},\ \text{and}\ \mathbf{R}_{{\bf{H}}_{p}{\bf{H}}_{p}}$. These covariance matrices enables the estimator to effectively account for channel correlations, thereby facilitating accurate denoising and interpolation across both time and frequency.
Despite its optimality under the Gaussian assumption, the MMSE estimator incurs high computational complexity due to the required matrix inversion in $(\mathbf{R}_{{\bf{H}}_{p}{\bf{H}}_{p}}+\sigma^{2}\mathbf{I})^{-1}$. To address this issue, a low-complexity approximation, termed the 1D frequency-domain MMSE (1D-MMSE) estimator, was proposed in [33]. In this method, the temporal-domain channel correlation is neglected, so that only the frequency-domain correlation is taken into account for channel estimation. The 1D-MMSE estimator is given as
@@
Despite their solid theoretical foundations, SP-based methods have several limitations. For instance, MMSE relies on prior knowledge of channel statistics, such as noise variance and channel correlation, which may be difficult to obtain or inaccurate in practical scenarios. Furthermore, MMSE estimation requires matrix inversion, incurring significant computational overhead per channel estimation. Also, the performance of these methods tends to degrade in highly dynamic or nonlinear channel environments, limiting their adaptability and robustness.
@@ IV-B1 Insights on Two-Stage Attention Encoder
Under the assumption of a wide-sense stationary uncorrelated scattering (WSSUS) environment [33], the second-order statistics of OFDM channels exhibit a separable correlation function across frequency and temporal domains. Formally, the channel correlation matrix between two channel taps separated by frequency index difference $\Delta n$ and time index difference $\Delta m$ is factorized as follows:
@@
where $\mathbf{h}$ denotes $\operatorname{vec}(\mathbf{H})\in\mathbb{C}^{NM}$, $\mathbf{R}_{f}\in\mathbb{C}^{N\times N}$ and $\mathbf{R}_{t}\in\mathbb{C}^{M\times M}$ represent the frequency and temporal domain correlation matrices, respectively, as visualized in Fig. 3. We note that the approximation ($\approx$) in (21) arises because the WSSUS assumption may not perfectly hold in practical scenarios. The representation of the full covariance matrix ${\bf{R}}_{\text{full}}$ as the Kronecker product ($\otimes$) between the frequency-domain correlation ${\bf{R}}_{f}$ and the temporal-domain correlation ${\bf{R}}_{t}$ (21) implies that the frequency-temporal domain correlation structure is separable. This allows for a compact and computationally efficient representation of the channel covariance matrix. Building on this insight, we propose the following two-stage encoder design.
@@
This two-stage encoder enables the model to effectively learn the joint frequency-temporal correlation structure, which is critical for the accurate MMSE estimation. Crucially, we note that this encoder design represents a key distinction of the proposed A-MMSE learning mechanism compared to conventional DNN-based methods. Specifically, existing approaches typically adopt standard learning architectures such as CNN [7] or Attention Transformer [13], without explicitly tailoring them to the statistical properties inherent in OFDM channels. In contrast, our method explicitly incorporates the inherent temporal and frequency-domain correlation structures of OFDM channels into the learning process. This integration not only enhances learning efficiency but also improves the model’s capability to generalize under practical channel conditions. In this sense, the proposed A-MMSE approach is distinguished not only by its design, which enables the learning of a linear estimation filter rather than directly estimating the OFDM channel matrix (as illustrated in Fig.1), but also by its advanced learning mechanism that explicitly reflects the inherent correlation structure of OFDM channels.
@@ IV-B2 Multi-Head Self-Attention Mechanism
Subsequently, the embedding matrix $\mathbf{X}_{\text{Emb}}$ is fed into the MHA module $\left(\mathcal{F}_{\text{MHA}}(\cdot)\right)$ with the number of heads fixed at $L$ (one head per each pilot). The operation of the MHA module $\mathcal{F}_{\text{MHA}}(\cdot)$ is described in (12)–(15). We note that this allows each head to learn distinct frequency domain correlation patterns. Following the MHA module, Add & Layer Normalization is applied, succeeded by a Feed-Forward Network (FFN), and subsequently another Add & Layer Normalization, in sequential order. Details on Add & Norm and FFN are provided in [11]. After these processes, the encoder extracts frequency-domain statistical information, which can be formally represented as
$\mathbf{Y}_{\text{Freq}}=\mathcal{E}_{\text{Frequency}}\left(\mathcal{F}_{\text{MHA}}(\mathbf{X}_{\text{Emb}}|\theta_{\text{MHA, Freq}})|\theta_{\text{Freq}}\right)$ (23)
@@
$\mathbf{Y}_{\text{Dec}}=\mathcal{F}_{\text{ResFC}}\left(\mathbf{Y}_{\text{Temp}}|\theta_{\text{ResFC}}\right),$ (26)

### v2

@@ Abstract
In orthogonal frequency division multiplexing (OFDM), accurate channel estimation is crucial. Classical signal processing based approaches, such as minimum mean-squared error (MMSE) estimation, often require second-order statistics that are difficult to obtain in practice. Recent deep neural networks based methods have been introduced to address this; yet they often suffer from high inference complexity. This paper proposes an Attention-aided MMSE (A-MMSE), a novel model-based DNN framework that learns the optimal MMSE filter via the Attention Transformer. Once trained, the A-MMSE estimates the channel through a single linear operation for channel estimation, eliminating nonlinear activations during inference and thus reducing computational complexity. To enhance the learning efficiency of the A-MMSE, we develop a two-stage Attention encoder, designed to effectively capture the channel correlation structure. Additionally, a rank-adaptive extension of the proposed A-MMSE allows flexible trade-offs between complexity and channel estimation accuracy. Extensive simulations with 3GPP TDL channel models demonstrate that the proposed A-MMSE consistently outperforms other baseline methods in terms of normalized MSE across a wide range of signal-to-noise ratio (SNR) conditions. In particular, the A-MMSE and its rank-adaptive extension establish a new frontier in the performance-complexity trade-off, providing a powerful yet highly efficient solution for practical channel estimation.
@@ I Introduction
Beyond traditional SP-based approaches, sensing-aided channel estimation has emerged within the integrated sensing and communication (ISAC) frameworks [7, 8]. These methods leverage communication signal echoes to extract sensing parameters, which then serve as prior information for channel refinement. Despite showing promise in exploiting communication-sensing synergy, these approaches require additional computational resources for sensing processing.
Recently, inspired by the remarkable success of deep neural networks (DNNs) across various engineering fields, DNN-aided channel estimation for OFDM has gained considerable attention. DNNs, as pre-trained neural network models, can effectively capture complex data relationships without explicit prior statistical knowledge. Leveraging this data-driven approach, several prior works have explored DNN-aided OFDM channel estimation strategies. In [9], an OFDM channel slot was treated as image data. Based on this, a super-resolution network and a denoising neural network were simultaneously employed to estimate the OFDM channel (i.e., ChannelNet). In [10], using a similar modeling approach to [9], ReEsNet [10] was proposed, where a residual learning-based DNN was used to reconstruct the OFDM channel. This method achieves lower computational cost and higher accuracy than [9]. In [11], a residual convolutional network combined with bilinear interpolation was proposed for OFDM channel estimation, offering enhanced flexibility to varying pilot patterns and reduced complexity. Beyond deterministic models, [12] optimized a deep generative model using compressive pilot measurements for high-dimensional channel estimation. This method leverages learned priors to reduce pilot overhead and surpass traditional sparse recovery techniques.
@@
Via numerical simulations across diverse 3GPP TDL scenarios, we show that the proposed A-MMSE and RA-A-MMSE significantly outperform traditional SP-based approaches (e.g., LS and MMSE) and recent DNN-based approaches (e.g., ChannelNet [9] and Channelformer [15]) in terms of channel estimation accuracy. In particular, the A-MMSE achieves approximately $43\%$ lower normalized MSE (NMSE) compared to conventional MMSE. We attribute these performance gains to the Attention Transformer’s ability to effectively capture nonstationary channel statistics, even under challenging high-mobility and frequency-selective fading conditions. Comparing to Channelformer, the A-MMSE achieves approximately $48\%$ lower NMSE. We note that these gain stem from our novel two-stage Attention encoder architecture, where frequency domain features are first extracted and subsequently refined through temporal correlation modeling. These channel estimation improvements directly translate to symbol detection performance gains, with A-MMSE achieving $13.1\%$ and $26.6\%$ lower bit error rate (BER) compared to MMSE and Channelformer respectively in the challenging mmWave scenario at $25$ dB SNR. Furthermore, the proposed method surpasses recent sensing-aided approaches [7], achieving up to a $35\%$ lower NMSE in high-mobility scenarios without the added complexity of echo signal processing. Moreover, the RA-A-MMSE further reduces computational complexity through adaptive rank reduction, maintaining over $99.6\%$ of the performance achieved by the full-rank A-MMSE. As a result, the RA-A-MMSE requires only $1.5\%$ of the computational complexity of Channelformer or MMSE (i.e., a $98.5\%$ reduction), while achieving approximately $36\%$ lower NMSE. The proposed methods also exhibit strong robustness to signal-to-ratio (SNR) mismatch, retaining nearly full performance ($\geq 98.1\%$) across a wide range of unseen SNR conditions. This underscores the practical efficiency and suitability of the proposed method for real-world 5G/6G implementations.
In summary, the proposed A-MMSE method offers several advantages over existing approaches. First, the proposed A-MMSE with novel two-stage Attention encoder effectively incorporates the OFDM channel correlation structure, thereby significantly improving channel estimation accuracy and robustness, achieving around $43\%$ and $48\%$ lower NMSE compared to MMSE and Channelformer, respectively. Second, once trained, the A-MMSE employs purely linear computations for OFDM channel estimation, significantly reducing computational complexity compared to conventional DNN-based methods that heavily rely on numerous nonlinear operations. Especially, the RA-A-MMSE dynamically adjusts its computational complexity according to available hardware resources, requiring $99.8\%$ fewer computations than the average of conventional DNN-based methods. This enables both efficient and flexible implementation in practical systems. Third, the A-MMSE also achieves up to $35\%$ lower NMSE than recent sensing-aided method in high-mobility scenarios, without the need for additional sensing signal processing.
@@
Notations: We denote transpose by $(\cdot)^{\mathsf{T}}$, Hermitian transpose by $(\cdot)^{\mathsf{H}}$, complex conjugate by $(\cdot)^{*}$, column‑wise vectorization by $\operatorname{vec}(\cdot)$, real and imaginary parts of a complex quantity by $\Re(\cdot)$ and $\Im(\cdot)$, the magnitude by $|\cdot|$, and the Euclidean and Frobenius norms by $\lVert\cdot\rVert_{2}$ and $\lVert\cdot\rVert_{F}$, respectively. Element‑wise (Hadamard) multiplication is denoted by $\circ$, the Kronecker product by $\otimes$, and statistical expectation is written as $\mathbb{E}[\cdot]$. The set difference is denoted by $\setminus$, horizontal (row-wise) stacking by $\cdot$$\cdot$], and vertical (column-wise) stacking by $\cdot$$\cdot$].
@@ III-A SP-based Approach
The SP-based OFDM channel estimation strategies have been extensively studied in [6].
Among the various approaches, this work intentionally focuses on pilot-assisted channel estimation, due to its robustness and practicality in wireless communication scenarios. In particular, the LS and MMSE estimators are commonly used as standard methods.
@@
The LS estimator requires no prior statistical knowledge of the channel or noise, making it simple and straightforward to implement. However, its performance may degrade significantly, especially at low SNR conditions or when the channel matrix $\bf{H}$} is ill-conditioned.
@@
Under the joint Gaussian assumption, the MMSE estimator is well-established to achieve optimal MSE performance. Unlike the LS estimator, it exploits knowledge of the covariance matrices $\mathbf{R}_{\operatorname{vec}(\mathbf{H}),\mathbf{H}p}$ and $\mathbf{R}_{{\bf{H}}_{p}{\bf{H}}_{p}}$, which enable effective denoising and interpolation by accounting for channel correlations across time and frequency.
Despite its optimality under the Gaussian assumption, the MMSE estimator incurs high computational complexity due to because of the high dimensionality of the covariance matrix in $\mathbf{R}_{\operatorname{vec}(\mathbf{H}){\bf{H}}_{p}}\in\mathbb{C}^{NM\times L}$. To address this issue, a low-complexity approximation, termed the 1D frequency-domain MMSE (1D-MMSE) estimator, is proposed. In this method, the temporal-domain channel correlation is neglected, so that only the frequency-domain correlation is taken into account for channel estimation. The 1D-MMSE estimator is given as
@@
Sensing-aided method: Beyond these purely pilot-based methods, recent research in ISAC has explored sensing-aided extensions that enhance traditional estimators by incorporating auxiliary information. The general framework can be expressed as
$\hat{\mathbf{H}}_{\text{sensing}}=g(\hat{\mathbf{H}}_{\text{init}},\theta),$ (11)
where $\hat{\mathbf{H}}_{\text{init}}$ is the initial estimate (e.g., from LS or MMSE) and $\theta$ represents auxiliary information from environmental parameters [7] and channel statistics [8]. Enhancement functions include additive refinement ($\hat{\mathbf{H}}=\hat{\mathbf{H}}_{\text{init}}+\Delta\mathbf{H}(\theta)$), multiplication refinement ($\hat{\mathbf{H}}=\mathbf{W}(\theta)\hat{\mathbf{H}}_{\text{init}}$), and projection-based methods. While these approaches can improve estimation accuracy, they require additional sensing infrastructure and computational resources.
@@
Despite their solid theoretical foundations, SP-based methods have several limitations. For instance, MMSE relies on prior knowledge of channel statistics, such as noise variance and channel correlation, which may be difficult to obtain or inaccurate in practical scenarios. Furthermore, MMSE estimation requires matrix inversion, incurring significant computational overhead per channel estimation. While sensing-aided extensions can potentially enhance estimation accuracy by incorporating auxiliary environmental information, they require additional sensing hardware and processing units, which may increase system complexity and cost. Such additional infrastructure could be less suitable for resource-constrained or cost-sensitive applications.
@@ IV-B1 Insights on Two-Stage Attention Encoder
Under the assumption of a wide-sense stationary uncorrelated scattering (WSSUS) environment and separable scattering function [35], the second-order statistics of OFDM channels exhibit a separable correlation function across frequency and temporal domains. Formally, the channel correlation matrix between two channel taps separated by frequency index difference $\Delta n$ and time index difference $\Delta m$ is factorized as follows:
@@
where $\mathbf{h}$ denotes $\operatorname{vec}(\mathbf{H})\in\mathbb{C}^{NM}$, $\mathbf{R}_{f}\in\mathbb{C}^{N\times N}$ and $\mathbf{R}_{t}\in\mathbb{C}^{M\times M}$ represent the frequency and temporal domain correlation matrices, respectively, as visualized in Fig. 3. A detailed proof of this Kronecker product structure is provided in Appendix A. We note that the approximation ($\approx$) in (22) arises because the WSSUS assumption may not perfectly hold in practical scenarios. The representation of the full covariance matrix ${\bf{R}}_{\text{full}}$ as the Kronecker product ($\otimes$) between the frequency-domain correlation ${\bf{R}}_{f}$ and the temporal-domain correlation ${\bf{R}}_{t}$ (22) implies that the frequency-temporal domain correlation structure is separable. This allows for a compact and computationally efficient representation of the channel covariance matrix. Building on this insight, we propose the following two-stage encoder design.
@@
This two-stage encoder enables the model to learn the joint frequency–temporal correlation structure; each stage uses a single encoder block, which is critical for the accurate MMSE estimation. Crucially, we note that this encoder design represents a key distinction of the proposed A-MMSE learning mechanism compared to conventional DNN-based methods. Specifically, existing approaches typically adopt standard learning architectures such as CNN [9] or Attention Transformer [15], without explicitly tailoring them to the statistical properties inherent in OFDM channels. In contrast, our method explicitly incorporates the inherent temporal and frequency-domain correlation structures of OFDM channels into the learning process. This integration not only enhances learning efficiency but also improves the model’s capability to generalize under practical channel conditions. In this sense, the proposed A-MMSE approach is distinguished not only by its design, which enables the learning of a linear estimation filter rather than directly estimating the OFDM channel matrix (as illustrated in Fig.1), but also by its advanced learning mechanism that explicitly reflects the inherent correlation structure of OFDM channels.
@@ IV-B2 Multi-Head Self-Attention Mechanism
Subsequently, the embedding matrix $\mathbf{X}_{\text{Emb}}$ is fed into the MHA module $\left(\mathcal{F}_{\text{MHA}}(\cdot)\right)$ with the number of heads fixed at $L$
, where $L$ equals the total number of DM-RS pilot symbol (one head per pilot). Each complete pilot is split into real and imaginary parts, resulting in the $2L$ column dimension of $\mathbf{X}_{\mathrm{Emb}}$. For example, with Type-A DM-RS and additional Position =1=1 (i.e., two DM-RS OFDM symbols) using a comb-2 pattern over 6 reousrce blocks (RBs), that has 72 subcarriers, there are $72/2=36$ pilots per DM-RS symbol, hence $L=36\times 2=72$ pilots in total. In general,
$L=\left(\frac{N_{\mathrm{RB}}\cdot N_{\mathrm{SC/RB}}}{2}\right)\times N_{\mathrm{DM-RS\,sym}},$ (24)
where $N_{\mathrm{RB}}$ is the number of resource blocks, $N_{\mathrm{SC/RB}}=12$ is the number of subcarriers per RB, and $N_{\mathrm{DM-RS\,sym}}$ denotes the number of OFDM symbols carrying DM-RS. The operation of the MHA module $\mathcal{F}_{\text{MHA}}(\cdot)$ is described in (13)–(16). We note that this allows each head to learn distinct frequency domain correlation patterns. Following the MHA module, Add & Layer Normalization is applied, succeeded by a FFN, and subsequently another Add & Layer Normalization, in sequential order. Details on Add & Norm and FFN are provided in [13]. After these processes, the encoder extracts frequency-domain statistical information, which can be formally represented as
$\mathbf{Y}_{\text{Freq}}=\mathcal{E}_{\text{Frequency}}\left(\mathcal{F}_{\text{MHA}}(\mathbf{X}_{\text{Emb}}|\theta_{\text{MHA, Freq}})|\theta_{\text{Freq}}\right)\$(25)
@@
$\mathbf{Y}_{\text{Dec}}=\mathcal{F}_{\text{ResFC}}\left(\mathbf{Y}_{\text{Temp}}|\theta_{\text{ResFC}}\right)\in\mathbb{R}^{2L\times NM},$ (28)